1. Given below are the demand schedule and supply schedule for the labour market for supervisors. Remember that demand for labour represents the employers’ demand for workers, while supply represents the workers’ willingness to work. Graph the demand and supply curve on one graph and determine equilibrium in this market. STATE the equilibrium. Label the graph properly.

Please state the wage and quantity that establishes equilibrium.

Daily Wage for Supervisors
Quantity Demanded
(000s)
Quantity Supplied
(000s)
$200 560,000 40,000
$225 475,000 65,000
$230 375,000 100,000
$270 300,000 125,000
$300 295,000 295,000
$325 200,000 350,000
$340 100,000 465,000
$365 61,000 575,000

2. Calculate the coefficient of price elasticity of demand if the daily wage goes from $230 to $270. Is elasticity at this level inelastic or elastic?
so.
1
Expert's answer
2019-09-30T09:17:50-0400

1. The equilibrium wage is $300 and quantity is 295,000 workers.

2. Ed = (300,000 - 375,000)/(270 - 230)×(270 + 230)/(300,000 + 375,000) = -75,000/40×500/675,000 = - 1.39, so the demand is elastic at this level.